7.Find the coordinate of the point of interception of the lines: 2x + y = 5 ; x + y = 3
a.(-1,2)
b.(1,2)
c.(1,-1)
d.(2,1)

8.The equation of a straight line at point (2,3) and slope 5 is given by
a.y=5x-3
b.y=5x-10
c.y=5x+3
d.y=3x-5
1
Expert's answer
2019-04-10T13:09:04-0400

7 2x+y=5, x+y=3 -> 2x+y-(x+y)=5-3 -> x=2, y=3-2=1. Var. d.

8 y-3=5(x-2) -> y=5x-7